# A Survey of Vibe Coding with Large Language Models

### v1

@@ 2.2 Preliminary
Applying reinforcement learning to code generation requires executable feedback signals, evolving from basic mechanisms to sophisticated training paradigms. Early approaches combine pretrained language models with deep RL using unit test feedback and critical sampling, achieving strong performance on competitive benchmarks [245]. Execution-based methods leverage PPO with compiler feedback for real-time refinement [246]. Advanced RL frameworks employ multi-granularity unit test feedback [247], break generation into curriculum subtasks with fine-grained optimization [248], use ranking-based alignment mechanisms [62], and leverage Group Relative Policy Optimization with compiler feedback to achieve competitive performance [249].
@@
Beyond supervised generation, autonomous agents tackle complete software engineering tasks through specialized architectures and multi-agent collaboration. Single-agent systems introduce custom agent-computer interfaces achieving strong benchmark performance [12], combine structure-aware code search with spectrum-based fault localization for cost-effective issue resolution [13], and demonstrate that simple hierarchical localization can outperform complex agents [14, 250]. Repository-level code generation integrates programming tools with agent strategies showing substantial improvements [99, 251], while iterative retrieval-generation approaches improve baselines on repository evaluation benchmarks [251]. Multi-agent frameworks employ specialized roles across the development process. These assign distinct programmer, test designer, and test executor agents achieving high pass rates with reduced token consumption [104], encode standardized operating procedures into assembly-line paradigms with role-based collaboration achieving near-perfect task completion [252], and implement chat chain architectures with dual-agent communication covering the full software development lifecycle at minimal cost [10]. Specialized frameworks replicate human programming cycles [127], investigate test-driven development principles showing consistent improvements with upfront test provision [253], and analyze self-repair mechanisms revealing modest performance gains bottlenecked by self-feedback abilities [110].
@@
Effective agent systems require mechanisms for tool use, function calling, and execution infrastructure to interact with external systems and APIs. Function calling frameworks teach language models to self-supervise tool use with simple APIs requiring minimal demonstrations [93], use teacher models for tool-related instruction data generation [254], and employ multi-task learning on granular tasks with large-scale API datasets [255]. Execution and interaction environments provide lightweight RL frameworks with safe execution for multiple programming languages [26], use executable code as unified action space outperforming alternative formats with specialized instruction datasets and fine-tuned models [256], and integrate Monte Carlo Tree Search with external feedback for deliberate problem-solving [257]. Optimization and deployment advances automatically identify parallelizable function calls achieving significant latency speedup and cost savings [258], enable edge deployment with compact models matching large model capabilities locally [259], and coordinate specialized agents for tool selection, execution, and calibration achieving higher success rates [150]. Retrieval, evaluation, and infrastructure contributions provide large-scale benchmarks showing retrieval from diverse datastores significantly improves performance [260], decompose tool use capability into evaluation aspects for API manipulation [261], develop multi-agent systems for code review automation achieving strong results on critical review tasks [150], enable customizable conversable agents programmed in natural or code language [150], and provide massive instruction-API pair datasets demonstrating cross-language transfer enabling open-source models to outperform proprietary alternatives on new API calls [262].
@@
An alternative to reinforcement learning lies in supervised fine-tuning and instruction tuning methods, which have become foundational for code model training. Instruction evolution approaches iteratively evolve code instructions achieving competitive performance [5], while self-instruction methods bootstrap instruction-following capabilities [44]. Fully transparent self-alignment pipelines enable training without proprietary distillation [263], and large-scale instruction datasets combine multiple instruction generation approaches [41]. Specialized tuning addresses security-centric generation improving code security substantially [264], optimization code generation [265], and code editing, debugging, and refinement through self-debugging approaches and iterative refinement without additional training [110, 105, 266]. Foundation models for code employ repository-level pretraining with extended context windows [267], train on diverse languages with fill-in-the-middle objectives and PII redaction [6], incorporate GitHub issues and documentation with repository-level training [268], and achieve competition-level performance using massive sampling and filtering strategies [269].
@@ 3.1 Definition of Vibe Coding
In this paper, we define Vibe Coding as an engineering methodology for software development grounded in large language models [18, 19, 20, 21, 270]. At its core, as shown in Figure 2, it represents a dynamic triadic relationship among human developers, software projects, and Coding Agents. In this paradigm, humans evolve from direct code authors into intent articulators, context curators, and quality arbiters. Projects extend beyond static code repositories to become multifaceted information spaces encompassing codebases, databases, and domain knowledge. Coding Agents, as intelligent executors, perform code generation, modification, and debugging under the dual guidance of human intent and project constraints.
@@
The collaboration among the three parties can be modeled as a Constrained Markov Decision Process (Constrained MDP) [34], wherein humans define the goal space and constraint boundaries, projects provide the state space and transition constraints, and Agents execute policies and state transitions:
@@ 3.2 Why Vibe Coding
Vibe Coding enables individual developers to deliver team-scale capabilities. Production applications traditionally require coordinating frontend, backend, database, security, DevOps, and QA specialists with substantial overhead. Coding agents provide diverse expertise across domains [10, 150, 271]. Developers focus on requirements while agents implement across stacks [272]. This reduces learning overhead through context engineering. Individual developers now implement cloud infrastructure and performance optimizations without formal training, benefiting resource-constrained entities by compressing prototypes from weeks to days.
@@
Vibe Coding aims to balance development velocity and code quality—historically conflicting objectives. Traditional workflows trade delivery speed against testing rigor [273]. Vibe Coding can improve both through autonomous iteration decoupled from human constraints. Agents sustain 24/7 operation: automated testing, refactoring, and performance profiling [274]. Automating mechanical tasks liberates cognitive resources for design and optimization [275, 13]. Agents enable exhaustive exploration through rapid iteration [201, 105] while human oversight validates architectural decisions.
@@
Vibe Coding democratizes development by lowering technical barriers. Traditional development required extensive programming knowledge before implementing ideas. Natural language becomes the primary creation interface [4, 5, 1]. Domain experts—medical practitioners, educators, designers—articulate needs without computer science education [35]. This diversifies innovation sources, materializing underrepresented perspectives [276, 90]. Economic impact manifests through creator economy expansion: domain experts monetize specialized tools without technical co-founders. This parallels previous democratization waves, representing software literacy’s evolution from specialized skill to broadly accessible capability [277].
@@ 4.1.1 Pre-training Code Corpora
Code LLMs require large-scale training data from diverse sources [35, 278]. These models rely on large-scale code corpora primarily sourced from open platforms like GitHub and Stack Overflow, with quality filtering based on metrics such as repository stars, documentation completeness, and community engagement [279].
Training datasets vary significantly in composition and curation strategies, with two primary approaches emerging: depth-focused strategies that concentrate on popular languages for quality, and breadth-focused strategies that span diverse languages for coverage [280, 281]. The Stack dataset uses the depth-focused approach, providing 3.1 TB of permissively licensed source code across 30 programming languages with careful attention to licensing and data provenance [36, 282, 37], while CodeParrot similarly trains exclusively on GitHub code to achieve depth in popular languages [278]. In contrast, breadth-focused approaches span diverse languages for maximum coverage. The Stack v2 expands coverage substantially to 67.5 TB across 619 programming and markup languages [37, 38], GPT-Neo utilizes diverse mixed corpora including "the Pile" [278], while CodeLlama employs SlimPajama’s 627 billion tokens paired with code from The Stack [282] and Arctic-SnowCoder uses a filtered combination yielding 400 billion tokens [37].
Data quality has driven sophisticated processing pipelines, with RefineCode incorporating over 130 carefully crafted language-specific filtering and cleaning rules across 607 languages to ensure syntactic and semantic validity [38], and SwallowCode implementing comprehensive four-stage pipelines performing syntax validation, quality assessment, deduplication, and LLM-based rewriting to enhance training data quality [39].
@@ 4.1.2 Instruction & Preference Datasets
Beyond raw code corpora, instruction-following requires curated training data. Accordingly, code LLMs have demonstrated exceptional capabilities across diverse programming tasks [279, 5], with instruction tuning as a critical technique for enhancing instruction-following abilities and solution generation quality [5]. These instruction datasets are typically compiled from two primary sources: permissively licensed source code repositories and synthetically constructed instructional data [279]. Notable instruction datasets in the field include CommitPack with its impressive 4 TB of raw data carefully filtered to 2GB of high-quality code paired with commit message instructions that capture developer intent [40], OpenCodeInstruct as the largest openly available dataset with 5 million diverse samples covering multiple programming languages and task types [41], and SynCode-Instruct providing nearly 20 billion tokens of synthetic instruction-following examples [42]. For preference learning applications, CodeUltraFeedback provides 10,000 coding instructions each paired with four distinct responses from 14 different LLMs, all evaluated comprehensively by GPT-3.5 to establish quality rankings [43], while the PLUM Framework innovatively employs GPT-4-generated unit tests to establish automated preference rankings based on functional correctness [283].
High-quality instruction dataset construction methodologies have evolved significantly from expensive and time-consuming human annotation processes toward more scalable synthetic generation approaches that maintain quality while reducing costs [41, 2]. The Self-Instruct paradigm pioneered the approach of bootstrapping from limited seed examples, using language models to generate new instructions through carefully designed instruction-input-output pipelines [44], while Evol-Instruct introduced the concept of progressive complexity increase by iteratively rewriting and enhancing initial instructions to create more challenging and diverse training examples [284]. OSS-Instruct advances the field by systematically integrating real-world open-source code snippets as seeds for generating diverse, contextually realistic instruction data that better reflects actual programming scenarios [285]. Quality enhancement techniques have become sophisticated, employing approaches like CoachLM’s LLM-based iterative rewriting to improve instruction clarity and task alignment, and CaR (Critique and Revise) approaches that use specialized scoring models aligned with expert preferences to filter and refine synthetic data, with the effectiveness of these synthetic data approaches demonstrated by systems like Nvidia’s Nemotron-4 achieving state-of-the-art performance using 98% synthetically generated alignment data [286].
@@ 4.2.1 Pre-training Objectives
Large language models for code leverage transfer learning paradigms inherited from BERT and GPT architectures [45, 287, 207, 47, 46], with pre-training on large-scale data before task-specific fine-tuning proving particularly effective where LLMs outperform domain-specific models in limited labeled data scenarios [47]. Code-specific pre-training objectives surpass traditional masked language modeling approaches, with innovations including GraphCodeBERT’s data flow prediction that captures semantic relationships in code [47, 48] and variable naming tasks in CodeT5 and DOBF that leverage programming language structure [47, 49, 7].
@@
Pre-training serves the fundamental purpose of instilling general-purpose knowledge and pattern recognition capabilities from large-scale unlabeled datasets [50, 51], with several well-established paradigms dominating the landscape. Autoregressive language modeling predicts the next token in a sequence based on all previous context [288, 51], while masked language modeling focuses on predicting randomly masked tokens using both preceding and following context for bidirectional understanding [288, 51]. Alternative objectives that have gained traction include infilling tasks that predict missing spans within sequences and prefix language modeling that generates continuations from given prefixes [50].
@@
Masked Language Modeling is the dominant approach for code understanding tasks, building directly on BERT’s remarkable success in natural language processing [289, 53, 45]. This approach is primarily adopted by encoder-only models like CodeBERT, which utilizes RoBERTa-based architectures for bidirectional code understanding [52]. The standard implementation samples approximately 15% of input tokens, with 80% of selected tokens replaced by a special [MASK] token to enable bidirectional context-based prediction [53].
@@
Implementation details vary across models, with CodeBERT applying MLM augmented with Replaced Token Detection to improve robustness [54, 290, 52], while GraphCodeBERT extends this with structure-aware features discussed below [48].
@@
Autoregressive language modeling serves as the foundational approach for code generation tasks by predicting the next token based solely on preceding context [289]. Decoder-only models like CodeGPT effectively leverage GPT architecture’s powerful Transformer-based decoders for this purpose [291, 45]. Recent models like CodeLlama have extended this paradigm through innovative fill-in-the-middle objectives that enable models to generate missing code segments given both prefix and suffix contexts, proving valuable for code completion tasks [292, 293].
@@
In contrast to unidirectional prediction, denoising objectives have proven highly effective for encoder-decoder architectures through span-level reconstruction tasks [53]. PLBART first applied denoising approaches to code by building on BART’s sequence-to-sequence architecture [55, 56, 290, 294], while modern unified models increasingly adopt sophisticated denoising strategies in comprehensive multi-task frameworks exemplified by CodeT5+ which synergistically combines span denoising with contrastive learning objectives [57, 203].
@@
Structure-aware objectives incorporate the unique structural characteristics inherent to programming languages, including abstract syntax trees, control flow graphs, and data flow relationships [52, 45]. GraphCodeBERT incorporates data flow graphs that explicitly encode semantic relationships between variables across code segments [47, 48], while introducing specialized structure-aware tasks such as code structure edge prediction that require understanding syntactic and semantic code organization [295, 48]. Contrastive learning explicitly minimizes representation distance between semantically similar code functions while simultaneously maximizing distance between dissimilar code, and is a powerful technique for training code models [296, 297]. UniXcoder uses this approach through sophisticated multi-modal contrastive learning that unifies natural language comments and linearized abstract syntax tree representations into a shared semantic space [53, 298], while ContraCode leverages automated compiler transformations to generate semantically equivalent code variations for robust contrastive data augmentation [296]. Multimodal pre-training strategies leverage multiple complementary data modalities to enhance model capabilities, with CodeBERT first learning general-purpose code representations through carefully aligned bimodal natural language-programming language pairs [52, 53, 295, 294], while advanced models like CodeT5+ implement sophisticated two-stage training procedures involving initial unimodal pre-training on massive code-only data to establish programming knowledge, followed by bimodal pre-training on carefully curated code-text pairs with explicit cross-modal alignment learning to bridge natural and programming languages [57, 279].
@@
SWE-Gym [299] 2025; ~2.4k tasks, including 230 lightweight instances Python; project-level bug fixing and editing tasks Collected from real pull requests in 11 open-source Python projects with executable environments and unit tests Contains complete code context, executable test environments, and unit tests; suitable for training RL coding agents Project-level, with repository context, directly executable
@@
Codeforces-CoTs [300] 2025; ~10k Codeforces problems, ~100k total samples, ~9.69GB C++ and Python competitive programming problems with chain-of-thought Collected 10k+ problems from Codeforces, using DeepSeek R1 to generate up to five reasoning trajectories per problem Problems accompanied by model-generated code and reasoning trajectories; ~84% Python solutions pass public tests Problem-level; includes multiple subsets (solutions, solutions_w_editorials, etc.)
@@
SWE-Fixer [301] 2025; 110k high-quality samples filtered from 2.3k repositories and 331k instances Multi-language, including code retrieval and editing Utilized GitHub events to crawl issue-PR pairs, limited to three files or fewer, removed parsing errors, generated chain-of-thought Each instance contains issue description, code context, and patches; suitable for supervised and RL training on retrieval/editing tasks File-level and project-level; includes chain-of-thought
@@
KodCode [302] 2024; 447k problem-solution-test triples Synthetic algorithmic/coding exercises Generated through multi-stage pipeline: problem synthesis, solution generation, unit test creation, and post-processing All problems equipped with verifiable unit tests, covering various difficulty levels; suitable for SFT and RL training Function-level; fully synthetic, emphasizes test executability
@@
Code-R1 [303] 2025; 12k RL samples (2k LeetCode + 10k filtered from TACO) Programming problems for RL training Selected from LeetCode and TACO verified instances with reliable tests, providing prompts and unit tests Emphasizes sandbox execution for reward verification, avoids test noise; directly usable for GRPO optimization Function-level; each contains prompt and tests
@@
Z1-Code [304] 2024; 107k coding problems Multi-language coding problems with both short and long reasoning trajectories Filtered from public problem sets and synthetic problems, annotated with short-chain and long-chain reasoning Provides two reasoning depths to improve model generalization and avoid over-thinking Problem-level; includes solutions and reasoning
@@
rStar-Coder [305] 2024; 418k competitive problems and 580k long-reasoning solutions Multi-language competitive programming problems Aggregated existing competitive problems and synthesized new ones, generated input-output tests and long-reasoning solutions Each problem has verification tests and in-depth reasoning solutions; suitable for training interpretable coding models Problem-level; includes tests and reasoning

### v2

@@ 2.2 Preliminary
Applying reinforcement learning to code generation requires executable feedback signals, evolving from basic mechanisms to sophisticated training paradigms. Early approaches combine pretrained language models with deep RL using unit test feedback and critical sampling, achieving strong performance on competitive benchmarks [le2022coderl]. Execution-based methods leverage PPO with compiler feedback for real-time refinement [shojaee2023execution]. Advanced RL frameworks employ multi-granularity unit test feedback [liu2023rltf], break generation into curriculum subtasks with fine-grained optimization [dou2024stepcoder], use ranking-based alignment mechanisms [shen2023pangu], and leverage Group Relative Policy Optimization with compiler feedback to achieve competitive performance [deepseekai2024deepseek].
@@
Beyond supervised generation, autonomous agents tackle complete software engineering tasks through specialized architectures and multi-agent collaboration. Single-agent systems introduce custom agent-computer interfaces achieving strong benchmark performance [yang2024swe], combine structure-aware code search with spectrum-based fault localization for cost-effective issue resolution [zhang2024autocoderover], and demonstrate that simple hierarchical localization can outperform complex agents [xia2024agentless, ma2025alibaba]. Repository-level code generation integrates programming tools with agent strategies showing substantial improvements [zhang2024codeagent, pan2024enhancing], while iterative retrieval-generation approaches improve baselines on repository evaluation benchmarks [pan2024enhancing]. Multi-agent frameworks employ specialized roles across the development process. These assign distinct programmer, test designer, and test executor agents achieving high pass rates with reduced token consumption [huang2023agentcoder], encode standardized operating procedures into assembly-line paradigms with role-based collaboration achieving near-perfect task completion [hong2024metagpt], and implement chat chain architectures with dual-agent communication covering the full software development lifecycle at minimal cost [qian2023chatdev]. Specialized frameworks replicate human programming cycles [islam2024mapcoder], investigate test-driven development principles showing consistent improvements with upfront test provision [mathews2024test], and analyze self-repair mechanisms revealing modest performance gains bottlenecked by self-feedback abilities [chen2023teaching].
@@
Effective agent systems require mechanisms for tool use, function calling, and execution infrastructure to interact with external systems and APIs. Function calling frameworks teach language models to self-supervise tool use with simple APIs requiring minimal demonstrations [schick2023toolformer], use teacher models for tool-related instruction data generation [yang2023gpt4tools], and employ multi-task learning on granular tasks with large-scale API datasets [abdelaziz2024granite]. Execution and interaction environments provide lightweight RL frameworks with safe execution for multiple programming languages [yang2023intercode], use executable code as unified action space outperforming alternative formats with specialized instruction datasets and fine-tuned models [wang2024executable], and integrate Monte Carlo Tree Search with external feedback for deliberate problem-solving [zhou2023language]. Optimization and deployment advances automatically identify parallelizable function calls achieving significant latency speedup and cost savings [kim2024llm], enable edge deployment with compact models matching large model capabilities locally [erdogan2024tinyagent], and coordinate specialized agents for tool selection, execution, and calibration achieving higher success rates [wu2023autogen]. Retrieval, evaluation, and infrastructure contributions provide large-scale benchmarks showing retrieval from diverse datastores significantly improves performance [wang2024coderag], decompose tool use capability into evaluation aspects for API manipulation [liu2024evaluation], develop multi-agent systems for code review automation achieving strong results on critical review tasks [wu2023autogen], enable customizable conversable agents programmed in natural or code language [wu2023autogen], and provide massive instruction-API pair datasets demonstrating cross-language transfer enabling open-source models to outperform proprietary alternatives on new API calls [guo2024api].
@@
An alternative to reinforcement learning lies in supervised fine-tuning and instruction tuning methods, which have become foundational for code model training. Instruction evolution approaches iteratively evolve code instructions achieving competitive performance [luo2023wizardcoder], while self-instruction methods bootstrap instruction-following capabilities [wang2022self]. Fully transparent self-alignment pipelines enable training without proprietary distillation [wei2024selfcodealign], and large-scale instruction datasets combine multiple instruction generation approaches [ahmad2025opencodeinstruct]. Specialized tuning addresses security-centric generation improving code security substantially [he2024instruction], optimization code generation [ma2024llamoco], and code editing, debugging, and refinement through self-debugging approaches and iterative refinement without additional training [chen2023teaching, madaan2023self, ma2024lingma]. Foundation models for code employ repository-level pretraining with extended context windows [guo2024deepseek], train on diverse languages with fill-in-the-middle objectives and PII redaction [li2023starcoder], incorporate GitHub issues and documentation with repository-level training [lozhkov2024starcoder], and achieve competition-level performance using massive sampling and filtering strategies [li2022competition].
@@ 3.1 Definition of Vibe Coding
In this paper, we define Vibe Coding as an engineering methodology for software development grounded in large language models [karpathy2025, ray2025review, sapkota2025vibe, horvat2025vibe, moore2025vibe]. At its core, as shown in Figure 2, it represents a dynamic triadic relationship among human developers, software projects, and Coding Agents. In this paradigm, humans evolve from direct code authors into intent articulators, context curators, and quality arbiters. Projects extend beyond static code repositories to become multifaceted information spaces encompassing codebases, databases, and domain knowledge. Coding Agents, as intelligent executors, perform code generation, modification, and debugging under the dual guidance of human intent and project constraints.
@@
The collaboration among the three parties can be modeled as a Constrained Markov Decision Process (Constrained MDP) [altman2021constrained], wherein humans define the goal space and constraint boundaries, projects provide the state space and transition constraints, and Agents execute policies and state transitions:
@@ 3.2 Why Vibe Coding
Vibe Coding enables individual developers to deliver team-scale capabilities. Production applications traditionally require coordinating frontend, backend, database, security, DevOps, and QA specialists with substantial overhead. Coding agents provide diverse expertise across domains [qian2023chatdev, wu2023autogen, dong2024villageragent]. Developers focus on requirements while agents implement across stacks [dong2023self]. This reduces learning overhead through context engineering. Individual developers now implement cloud infrastructure and performance optimizations without formal training, benefiting resource-constrained entities by compressing prototypes from weeks to days.
@@
Vibe Coding aims to balance development velocity and code quality—historically conflicting objectives. Traditional workflows trade delivery speed against testing rigor [dima2018waterfall]. Vibe Coding can improve both through autonomous iteration decoupled from human constraints. Agents sustain 24/7 operation: automated testing, refactoring, and performance profiling [Janzen2005TestdrivenDC]. Automating mechanical tasks liberates cognitive resources for design and optimization [goues2019automated, zhang2024autocoderover]. Agents enable exhaustive exploration through rapid iteration [shinn2023reflexion, madaan2023self] while human oversight validates architectural decisions.
@@
Vibe Coding democratizes development by lowering technical barriers. Traditional development required extensive programming knowledge before implementing ideas. Natural language becomes the primary creation interface [nijkamp2023codegen, luo2023wizardcoder, brown2020language]. Domain experts—medical practitioners, educators, designers—articulate needs without computer science education [chen2021evaluating]. This diversifies innovation sources, materializing underrepresented perspectives [xi2023rise, park2023generative]. Economic impact manifests through creator economy expansion: domain experts monetize specialized tools without technical co-founders. This parallels previous democratization waves, representing software literacy’s evolution from specialized skill to broadly accessible capability [clyburnesherin2019computational].
@@ 4.1.1 Pre-training Code Corpora
Code LLMs require large-scale training data from diverse sources [chen2021evaluating, xu2022systematic]. These models rely on large-scale code corpora primarily sourced from open platforms like GitHub and Stack Overflow, with quality filtering based on metrics such as repository stars, documentation completeness, and community engagement [jiang2024survey].
Training datasets vary significantly in composition and curation strategies, with two primary approaches emerging: depth-focused strategies that concentrate on popular languages for quality, and breadth-focused strategies that span diverse languages for coverage [albalak2024survey, longpre2023flan]. The Stack dataset uses the depth-focused approach, providing 3.1 TB of permissively licensed source code across 30 programming languages with careful attention to licensing and data provenance [wang2023mathpile, aryabumi2024code, wei2024arctic], while CodeParrot similarly trains exclusively on GitHub code to achieve depth in popular languages [xu2022systematic]. In contrast, breadth-focused approaches span diverse languages for maximum coverage. The Stack v2 expands coverage substantially to 67.5 TB across 619 programming and markup languages [wei2024arctic, huang2024opencoder], GPT-Neo utilizes diverse mixed corpora including "the Pile" [xu2022systematic], while CodeLlama employs SlimPajama’s 627 billion tokens paired with code from The Stack [aryabumi2024code] and Arctic-SnowCoder uses a filtered combination yielding 400 billion tokens [wei2024arctic].
Data quality has driven sophisticated processing pipelines, with RefineCode incorporating over 130 carefully crafted language-specific filtering and cleaning rules across 607 languages to ensure syntactic and semantic validity [huang2024opencoder], and SwallowCode implementing comprehensive four-stage pipelines performing syntax validation, quality assessment, deduplication, and LLM-based rewriting to enhance training data quality [fujii2025rewriting].
@@ 4.1.2 Instruction & Preference Datasets
Beyond raw code corpora, instruction-following requires curated training data. Accordingly, code LLMs have demonstrated exceptional capabilities across diverse programming tasks [jiang2024survey, luo2023wizardcoder], with instruction tuning as a critical technique for enhancing instruction-following abilities and solution generation quality [luo2023wizardcoder]. These instruction datasets are typically compiled from two primary sources: permissively licensed source code repositories and synthetically constructed instructional data [jiang2024survey]. Notable instruction datasets in the field include CommitPack with its impressive 4 TB of raw data carefully filtered to 2GB of high-quality code paired with commit message instructions that capture developer intent [muennighoff2023octopack], OpenCodeInstruct as the largest openly available dataset with 5 million diverse samples covering multiple programming languages and task types [ahmad2025opencodeinstruct], and SynCode-Instruct providing nearly 20 billion tokens of synthetic instruction-following examples [yang2024evaluating]. For preference learning applications, CodeUltraFeedback provides 10,000 coding instructions each paired with four distinct responses from 14 different LLMs, all evaluated comprehensively by GPT-3.5 to establish quality rankings [weyssow2024codeultrafeedback], while the PLUM Framework innovatively employs GPT-4-generated unit tests to establish automated preference rankings based on functional correctness [chen2024mastering].
High-quality instruction dataset construction methodologies have evolved significantly from expensive and time-consuming human annotation processes toward more scalable synthetic generation approaches that maintain quality while reducing costs [ahmad2025opencodeinstruct, ouyang2022training]. The Self-Instruct paradigm pioneered the approach of bootstrapping from limited seed examples, using language models to generate new instructions through carefully designed instruction-input-output pipelines [wang2022self], while Evol-Instruct introduced the concept of progressive complexity increase by iteratively rewriting and enhancing initial instructions to create more challenging and diverse training examples [xu2023wizardlm]. OSS-Instruct advances the field by systematically integrating real-world open-source code snippets as seeds for generating diverse, contextually realistic instruction data that better reflects actual programming scenarios [wei2023magicoder]. Quality enhancement techniques have become sophisticated, employing approaches like CoachLM’s LLM-based iterative rewriting to improve instruction clarity and task alignment, and CaR (Critique and Revise) approaches that use specialized scoring models aligned with expert preferences to filter and refine synthetic data, with the effectiveness of these synthetic data approaches demonstrated by systems like Nvidia’s Nemotron-4 achieving state-of-the-art performance using 98% synthetically generated alignment data [zheng2024opencodeinterpreter].
@@ 4.2.1 Pre-training Objectives
Large language models for code leverage transfer learning paradigms inherited from BERT and GPT architectures [devlin2019bert, qiu2020pre, zhou2024comprehensive, chirkova2023codebpe, raffel2019exploring], with pre-training on large-scale data before task-specific fine-tuning proving particularly effective where LLMs outperform domain-specific models in limited labeled data scenarios [chirkova2023codebpe]. Code-specific pre-training objectives surpass traditional masked language modeling approaches, with innovations including GraphCodeBERT’s data flow prediction that captures semantic relationships in code [chirkova2023codebpe, guo2020graphcodebert] and variable naming tasks in CodeT5 and DOBF that leverage programming language structure [chirkova2023codebpe, wang2021identifier, roziere2023code].
@@
Pre-training serves the fundamental purpose of instilling general-purpose knowledge and pattern recognition capabilities from large-scale unlabeled datasets [blouir2024birdie, jern2025agent], with several well-established paradigms dominating the landscape. Autoregressive language modeling predicts the next token in a sequence based on all previous context [li2025generated, jern2025agent], while masked language modeling focuses on predicting randomly masked tokens using both preceding and following context for bidirectional understanding [li2025generated, jern2025agent]. Alternative objectives that have gained traction include infilling tasks that predict missing spans within sequences and prefix language modeling that generates continuations from given prefixes [blouir2024birdie].
@@
Masked Language Modeling is the dominant approach for code understanding tasks, building directly on BERT’s remarkable success in natural language processing [liu2020multi, guo2022unixcoder, devlin2019bert]. This approach is primarily adopted by encoder-only models like CodeBERT, which utilizes RoBERTa-based architectures for bidirectional code understanding [feng2020codebert]. The standard implementation samples approximately 15% of input tokens, with 80% of selected tokens replaced by a special [MASK] token to enable bidirectional context-based prediction [guo2022unixcoder].
@@
Implementation details vary across models, with CodeBERT applying MLM augmented with Replaced Token Detection to improve robustness [huang2023program, yang2023enhancing, feng2020codebert], while GraphCodeBERT extends this with structure-aware features discussed below [guo2020graphcodebert].
@@
Autoregressive language modeling serves as the foundational approach for code generation tasks by predicting the next token based solely on preceding context [liu2020multi]. Decoder-only models like CodeGPT effectively leverage GPT architecture’s powerful Transformer-based decoders for this purpose [liu2022automatically, devlin2019bert]. Recent models like CodeLlama have extended this paradigm through innovative fill-in-the-middle objectives that enable models to generate missing code segments given both prefix and suffix contexts, proving valuable for code completion tasks [saad2025senai, li2025revisiting].
@@
In contrast to unidirectional prediction, denoising objectives have proven highly effective for encoder-decoder architectures through span-level reconstruction tasks [guo2022unixcoder]. PLBART first applied denoising approaches to code by building on BART’s sequence-to-sequence architecture [chakraborty2022natgen, ahmad2021unified, yang2023enhancing, lyu2024automatic], while modern unified models increasingly adopt sophisticated denoising strategies in comprehensive multi-task frameworks exemplified by CodeT5+ which synergistically combines span denoising with contrastive learning objectives [wang2023open, zheng2023survey].
@@
Structure-aware objectives incorporate the unique structural characteristics inherent to programming languages, including abstract syntax trees, control flow graphs, and data flow relationships [feng2020codebert, devlin2019bert]. GraphCodeBERT incorporates data flow graphs that explicitly encode semantic relationships between variables across code segments [chirkova2023codebpe, guo2020graphcodebert], while introducing specialized structure-aware tasks such as code structure edge prediction that require understanding syntactic and semantic code organization [lin2024scaling, guo2020graphcodebert]. Contrastive learning explicitly minimizes representation distance between semantically similar code functions while simultaneously maximizing distance between dissimilar code, and is a powerful technique for training code models [jain2020contrastive, zhou2024large]. UniXcoder uses this approach through sophisticated multi-modal contrastive learning that unifies natural language comments and linearized abstract syntax tree representations into a shared semantic space [guo2022unixcoder, dharma2023neural], while ContraCode leverages automated compiler transformations to generate semantically equivalent code variations for robust contrastive data augmentation [jain2020contrastive]. Multimodal pre-training strategies leverage multiple complementary data modalities to enhance model capabilities, with CodeBERT first learning general-purpose code representations through carefully aligned bimodal natural language-programming language pairs [feng2020codebert, guo2022unixcoder, lin2024scaling, lyu2024automatic], while advanced models like CodeT5+ implement sophisticated two-stage training procedures involving initial unimodal pre-training on massive code-only data to establish programming knowledge, followed by bimodal pre-training on carefully curated code-text pairs with explicit cross-modal alignment learning to bridge natural and programming languages [wang2023open, jiang2024survey].
@@
SWE-Gym [pan2024training] 2025; ~2.4k tasks, including 230 lightweight instances Python; project-level bug fixing and editing tasks Collected from real pull requests in 11 open-source Python projects with executable environments and unit tests Contains complete code context, executable test environments, and unit tests; suitable for training RL coding agents Project-level, with repository context, directly executable
@@
Codeforces-CoTs [penedo2025codeforces] 2025; ~10k Codeforces problems, ~100k total samples, ~9.69GB C++ and Python competitive programming problems with chain-of-thought Collected 10k+ problems from Codeforces, using DeepSeek R1 to generate up to five reasoning trajectories per problem Problems accompanied by model-generated code and reasoning trajectories; ~84% Python solutions pass public tests Problem-level; includes multiple subsets (solutions, solutions_w_editorials, etc.)
@@
SWE-Fixer [xie2025swe] 2025; 110k high-quality samples filtered from 2.3k repositories and 331k instances Multi-language, including code retrieval and editing Utilized GitHub events to crawl issue-PR pairs, limited to three files or fewer, removed parsing errors, generated chain-of-thought Each instance contains issue description, code context, and patches; suitable for supervised and RL training on retrieval/editing tasks File-level and project-level; includes chain-of-thought
@@
KodCode [xu2025kodcode] 2024; 447k problem-solution-test triples Synthetic algorithmic/coding exercises Generated through multi-stage pipeline: problem synthesis, solution generation, unit test creation, and post-processing All problems equipped with verifiable unit tests, covering various difficulty levels; suitable for SFT and RL training Function-level; fully synthetic, emphasizes test executability
@@
Code-R1 [code-r1] 2025; 12k RL samples (2k LeetCode + 10k filtered from TACO) Programming problems for RL training Selected from LeetCode and TACO verified instances with reliable tests, providing prompts and unit tests Emphasizes sandbox execution for reward verification, avoids test noise; directly usable for GRPO optimization Function-level; each contains prompt and tests
@@
Z1-Code [yu2025z1] 2024; 107k coding problems Multi-language coding problems with both short and long reasoning trajectories Filtered from public problem sets and synthetic problems, annotated with short-chain and long-chain reasoning Provides two reasoning depths to improve model generalization and avoid over-thinking Problem-level; includes solutions and reasoning
@@
rStar-Coder [liu2025rstar] 2024; 418k competitive problems and 580k long-reasoning solutions Multi-language competitive programming problems Aggregated existing competitive problems and synthesized new ones, generated input-output tests and long-reasoning solutions Each problem has verification tests and in-depth reasoning solutions; suitable for training interpretable coding models Problem-level; includes tests and reasoning